shausaesn wight -59q7 N, 4 0330, push ups.Fiad te nxmal sorce auui the Plooc on eadn hand and eacn Soot,assuming thot person hold this posihon. Li -0.3 31m ○ .475
Answer #1

The normal forces at both the hands and both the feet are different because the center of mass is at different distances from the hands and the feet. Let FNH and FNF be the normal forces exerted on the hands and on the feet, respectively.

Since, the body of the person is stationary at the moment shown in the figure,

FNF + FNH = W                 (1)

Also, the torques due to normal forces on hands will be equal and opposite to the torque due to the weight of the person at the center of mass (relative to the feet). So,

FNH(0.833 + 0.396) = W(0.833)

FNH = 497.301/1.229

FNH = 404.64 N

And from equation (1),

FNF = W - FNH = 192.36 N

Therefore, the normal force on each hand is 202.32 N and the normal force on each foot is 96.18 N.